OpenStudy (x__hazel__x):

Spencer rolled a number cube 50 times and recorded the results in this table. Outcome Number 1 8 2 10 3 7 4 7 5 9 6 9 What is the theoretical probability of the cube landing on a 2? What is the experimental probability of the cube landing on a 2? A. Theoretical - 16.7% Experimental - 20% B. Theoretical - 20% Experimental - 16.7% C. Theoretical - 10% Experimental - 20% D. Theoretical - 16.7% Experimental - 16.7%

OpenStudy (mathmale):

Theoretically, the six possible outcomes of the die throw have equal probability: 1 in 6, or 16.7%. Out of a total of 50 throws in an actual experiment, the number '2' showed up how many times? How would you use this figure and the total number of throws, 50, to calculate the experimental (actual) probability?

OpenStudy (x__hazel__x):

2 popped up 500 times (I think), and I am not sure for the last question. :3 @mathmale

OpenStudy (mathmale):

On what do you base your statement that "2 popped up 500 times?" The data table doesn't show that. Rather, it shows that 2 popped up 10 times OUT OF 50. This was your experimental result. What would be the theoretical result? As I said earlier, "Theoretically, the six possible outcomes of the die throw have equal probability."

OpenStudy (x__hazel__x):

Sorry, Im back!

OpenStudy (x__hazel__x):

the theoretical result would be 2 * 10 = 20

OpenStudy (x__hazel__x):

I mean, experimental

OpenStudy (mathmale):

You are calculating probabilities. Probabilities are always between 0 and 1, so '20' could not possibly be correct. Toss out the multiplication and try division instead. I see you typed '2' (which is an outcome, like 1, 3, 4, 5 and 6) TIMES 10, the number of times that 2 showed up. Again, this produces 20, which is not at all a probability. How many times was the number cube rolled? How many times did '2' come up? Find the experimental probability. Important: do not multiply here; instead, divide.

OpenStudy (x__hazel__x):

The number cube was rolled 50 times. 2 came up 10 times. The experimental probability is 1/5, but I am really not sure...

OpenStudy (mathmale):

You correctly divided 10 (the number of times that '2' came up) by 50 (the number of tosses of the cube). Result: 10/50, or 1/5, or .20, or 20%. What was the theoretical probability of obtaining a '2' on one toss?

OpenStudy (x__hazel__x):

0.2 = 2/10 = 20%?

OpenStudy (mathmale):

We've already discussed the experimental probability; it is 0.2 = 2/10 = 20%. I asked you to find the theoretical probability. Please do that. No need to spend much time on this, because we've already discussed it (above.) I wanted to make certain that you undrstand the difference between experimental and theoretical probabilities.

OpenStudy (x__hazel__x):

No, I dont think I do

OpenStudy (mathmale):

Nothing is perfect. But suppose you had a perfect die. Its sides are numbered {1, 2, 3, 4, 5, 6). Each of these 6 numbers has exactly the same chance of showing up. How many possible outcomes are there here to one toss of the die?

OpenStudy (x__hazel__x):

6?

OpenStudy (mathmale):

Yes. And how many of those 6 numbers are '2'??

OpenStudy (x__hazel__x):

1/6?

OpenStudy (mathmale):

Well, yes. But I asked how many times '2' shows up on this perfect die. How many?

OpenStudy (x__hazel__x):

Once?

OpenStudy (mathmale):

Right. So THAT is why the THEORETICAL probabililty of obtaining a '2' on one toss of our imaginary "perfect" die is 1/6. But now the plot thickens!! We take an imperfect die and toss it 50 times. We find that '2' shows up 10 times out of 50. This is the actual result of an experiment using a probably imperfect die. What is the EXPERIMENTAL probability of obtaining a '2'?

OpenStudy (mathmale):

Be sure to read our previous conversation over before answering this latest question. I am asking you about something we have already discussed.

OpenStudy (x__hazel__x):

How would we find the experimental? Im so sorry that this is taking forever.. Im a really slow person -.-

OpenStudy (mathmale):

Be sure to read our previous conversation over before answering this latest question. I am asking you about something we have already discussed. Different people learn at different rates; probability was once very difficult for me.

OpenStudy (x__hazel__x):

Me as well, and I will read it right now!

OpenStudy (x__hazel__x):

Okay, so the experimental is 20%

OpenStudy (mathmale):

Yes, the experimental prob. of obt. a '2' is 0.20. See the difference? Theoretical: what might happen in a perfect world, with a perfect die. Experimental: what we actually observe in an experiment we do ourselves. You toss the die 50 times; you count the number of 2s that come up, you divide that number by the number of tosses (50), and lo and behold, you have your experimental probability.

OpenStudy (x__hazel__x):

Ohhh okay! Thank you!

OpenStudy (mathmale):

My pleasure. Take care, Hazel, and good night.

OpenStudy (x__hazel__x):

Goodnight and thank you sooo much!